A garden, measuring 10m x 16m, is enclosed by a side walk of width 1m .the area of the side walk, in m^2 is
a) 8 b) 28 c) 56 d) 208 e) 216

Found 2 solutions by checkley77, nyc_function:
Answer by checkley77(12844)

10*16=160
(10+2)(16+2)
12*18=216
216-160=56 ans. C).

Answer by nyc_function(2741)

Here is the set up leading to a quadratic equation and then you can do the math.
160 = (10 + 2m)(16 + 2m)
Can you solve for m now?